I need to find the partial sum of 1+5+9+...+109
I understand how to find partial sum. I just haven't been able to figure out the common difference between the numbers.
So far, I have the equation set up as 109=1+(n-1)d
I just need to find d.
Could you explain to me how to find this answer?

Answer by MathTherapy(10552)


I need to find the partial sum of 1+5+9+...+109
I understand how to find partial sum. I just haven't been able to figure out the common difference between the numbers.
So far, I have the equation set up as 109=1+(n-1)d
I just need to find d.
Could you explain to me how to find this answer?

d or common difference is 4. Isn't 5 - 1 = 4?

Doesn't 9 - 5 = 4?